two charges of value 2 microcoulomb and -50 microcoulomb are placed 80cm apart calculate the distanceof the point from the smaller charge where the intensity is zero.​

  • two charges of value 2 microcoulomb and -50 microcoulomb are placed 80cm apart calculate the distanceof the point from the smaller charge where the intensity is zero.​

Grade:11

1 Answers

Shivam
3 years ago
80cm =0.8m 
E=E+ E- ,    E=0
E+=E-
K×2/x²= k×-50/(x-0.8)²
2/x²= -50/(x-0.8)²
1/x²=-25/(x-0.8)²
x-0.8= 5x
4x=0.8
x=0.2